OpenStudy (anonymous):

How would I go by dividing Functions? Ex: f(x)=4x-2 and g(x)-2x+1 F(-2)/g(1)

OpenStudy (anonymous):

plug -2 into your f function and 1 into your g function

OpenStudy (anonymous):

f(-2)=4(-2)-2=...? g(1)=-2(1)+1=...? then solve f(-2)/g(1)